# GAP-LA: GPU-Accelerated Performance-Driven Layer Assignment

### v1

@@ III-D Layer Assignment with Look Ahead Strategy
IV Experimental Results
@@
We implemented our GPU-accelerated layer assignment algorithm using C++ and CUDA on top of HeLEM-GR [25] and HeteroSTA [29, 30, 31], and evaluated the results on benchmarks from the ISPD25 contest [26] released by NVIDIA. The ISPD25 contest released 12 industrial designs synthesized with the Nangate 45nm open library technology node and specified an official evaluation flow to assess routing solution performance, as detailed in II. All experiments were conducted on a Linux server equipped with a 32-core Intel Xeon Gold CPU@2.90GHz and an NVIDIA A100 GPU. To match the contest evaluation setup, we used the official evaluator for overflow assessment and OpenROAD for evaluating TNS, WNS, and power consumption.
@@ IV-A Main Comparison with Top-3 Winners in the Contest
We obtained the top-3 contest winners’ results and reproduced their experiments on our hardware for comparison. Table II presents a comparison of quality scores and runtime against the ISPD 2025 contest’s top-3 winners. Our framework achieved the highest quality score in 9 out of 12 benchmark cases, with comparable performance in the remaining 3 cases. While our static timing analysis engine incurs a 25% runtime overhead due to parsing netlist (.v) and Synopsys design constraints (.sdc) files, the optimization method remains efficient and effective—demonstrating competitive performance relative to the top-3 winners.
@@
Table III presents the WNS, TNS, power consumption, and congestion results of four routers across 12 benchmark cases. The 2nd-place winner produces solutions with favorable performance and low congestion, while the 3rd-place winner demonstrates better performance at the expense of higher congestion. The 1st-place winner achieves a more balanced trade-off among metrics [26]. Our method delivers superior performance with acceptable congestion increase, which can be attributed to the critical net analysis-based net ordering and batching strategy. As shown in Table III, compared to the 1st-place winner, our router achieves 9.9% better WNS, 4.5% better TNS, 1.1% lower power consumption, and 3.8% higher congestion. Versus the 2nd-place winner, we achieve 0.3% better WNS, 5.4% better TNS, 1.5% lower power consumption, and 15.5% higher congestion. In contrast, the 3rd-place winner shows only 0.4% lower power consumption but suffers 4.4% worse WNS, 2.0% worse TNS, and 35.2% higher congestion compared to our approach.
IV-B Ablation Studies
| Case | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 | 11 | 12 | Avg. | |
@@ IV-B Ablation Studies
Table IV shows that our framework achieves better WNS results in 11 of the 12 ISPD2025 benchmark cases and better TNS results in all 12 cases. The net ordering and batching strategy enables more precise optimization of critical nets, significantly outperforming the naive strategy that ignores timing analysis information. This advantage is particularly evident in cases sensitive to critical net delay and capacitance, such as case 9. The look-ahead strategy also proves crucial in certain cases, including case 4.
@@
Table V presents the number of critical paths and nets for several selected cases. The results show that only a small fraction of nets (less than 1% of all nets) in the netlist play a significant role in timing closure. In case 8, focusing optimization efforts on the 9 critical nets with high timing cost weights is sufficient to achieve zero slack, demonstrating the effectiveness of our critical net selection strategy.
@@
In this paper, we present a GPU-accelerated performance-driven layer assignment framework that generates high-quality routing solutions from 2D routing results. The framework integrates a static timing analysis engine to schedule optimal net ordering and task batching for effective performance optimization. To address the delay cost optimization of large-scale designs with up to 10 million nets, we design GPU kernels for layer assignment using a look-ahead strategy. Compared with the top-3 winners of the contest, our router achieves 0.3%–9.9% better WNS and 2.0%–5.4% better TNS with acceptable runtime overhead.

### v2

@@ IV Experimental Results
| Quality Score | ||||
| 1st + GAP-LA† | 2nd + GAP-LA† | 3rd + GAP-LA† | [33] + GAP-LA† | |
| case1 | -1.161 (+2.058) | -1.223 (+1.863) | -1.182 (+1.276) | -1.189 (+3.829) |
| case2 | -0.054 (+0.616) | -0.127 (+0.849) | -0.196 (+0.290) | 0.013 (+1.016) |
| case3 | -1.996 (+0.001) | -1.965 (-0.112) | -2.023 (+0.199) | -2.032 (+9.747) |
| case4 | -0.369 (-0.036) | -0.453 (+0.854) | -0.435 (+1.805) | -0.319 (+1.805) |
| case5 | -3.808 (+1.779) | -6.167 (+4.872) | -4.456 (+0.857) | -6.015 (+6.555) |
| case6 | 0.384 (+0.126) | 0.366 (+0.033) | 0.329 (+0.385) | 0.397 (+0.711) |
| case7 | 1.532 (+0.260) | 1.535 (+0.124) | 1.524 (+0.007) | 1.551 (+0.193) |
| case8 | 0.426 (+0.423) | 0.446 (-0.031) | 0.425 (+0.263) | 0.491 (+0.741) |
| case9 | 1.347 (+0.084) | 1.588 (+0.073) | 1.410 (-0.018) | 1.379 (+9341.530) |
| case10 | 0.193 (+0.151) | 0.097 (+1.978) | 0.046 (+1.758) | 0.131 (+3.360) |
| case11 | 1.337 (-0.005) | 1.178 (+0.054) | 1.182 (+1.933) | 1.254 (+1.310) |
| case12 | 1.772 (+0.555) | 1.734 (+0.636) | 1.630 (+1.047) | 1.720 (+1.583) |
| Avg. Imp. | +0.501 | +0.933 | +0.817 | +781.032 |
| WNS | TNS | |||||||
| 1st + GAP-LA | 2nd + GAP-LA | 3rd + GAP-LA | [33] + GAP-LA | 1st + GAP-LA | 2nd + GAP-LA | 3rd + GAP-LA | [33] + GAP-LA | |
| case1 | -0.366 (+15.4%) | -0.366 (+8.2%) | -0.367 (+3.6%) | -0.364 (+31.8%) | -995.9 (+24.9%) | -986.8 (+26.4%) | -987.7 (+13.2%) | -992.5 (+32.7%) |
| case2 | -0.391 (+7.06%) | -0.387 (+8.7%) | -0.384 (+5.6%) | -0.389 (+11.2%) | -9594.3 (+7.4%) | -9541.2 (+11.4%) | -9478.6 (-1.4%) | -9635.2 (+11.8%) |
| case3 | -57.589 (-1.2%) | -56.475 (-1.1%) | -57.642 (-3.1%) | -56.905 (+63.9%) | -520231.3 (+0.2%) | -518380.2 (-0.7%) | -519322.0 (-0.9%) | -517720.5 (+40.2%) |
| case4 | -0.383 (-1.3%) | -0.384 (+21.5%) | -0.387 (+25.5%) | -0.390 (+45.1%) | -1678.2 (-4.2%) | -1673.6 (+24.8%) | -1669.5(+33.9%) | -1735.9 (+52.8%) |
| case5 | -0.276 (+9.7%) | -0.258 (+25.6%) | -0.253 (+8.2%) | -0.314 (+65.1%) | -17315.2 (+15.4%) | -17153.2 (+21.0%) | -16941.9 (+12.4%) | -17674.9 (+64.5%) |
| case6 | -0.249 (+6.4%) | -0.246 (+21.8%) | -0.251 (+9.0%) | -0.248 (+68.3%) | -52866.6 (+11.5%) | -52649.0 (+15.7%) | -52342.4 (+12.1%) | -52505.1 (+37.5%) |
| case7 | -0.643 (+63.4%) | -0.610 (+17.2%) | -0.614 (+3.0%) | -0.668 (+59.6%) | -111.0 (+82.9%) | -101.3 (+32.2%) | -102.0 (+37.2%) | -117.4 (+78.2%) |
| case8 | 0 (+100.0%) | 0 (-) | 0 (+100.0%) | 0 (+100%) | 0 (+100.0%) | 0 (-) | 0 (+100.0%) | 0 (+100.0%) |
| case9 | 0 (-) | 0 (-) | 0 (-) | 0 (+100%) | 0 (-) | 0 (-) | 0 (-) | 0 (+100.0%) |
| case10 | -0.348 (-1.6%) | -0.329 (+30.3%) | -0.333 (+24.4%) | -0.329 (+45.0%) | -1315.0 (+3.0%) | -1303.2 (+32.1%) | -1289.2 (+28.6%) | -1299.0 (+48.9%) |
| case11 | -0.403 (+7.0%) | -0.352 (+23.4%) | -0.351 (+5.8%) | -0.366 (+73.7%) | -27048.8 (+15.5%) | -25870.8 (+25.8%) | -25364.8 (+5.4%) | -26550.1 (+54.2%) |
| case12 | -0.193 (+13.1%) | -0.209 (+23.4%) | -0.194 (+5.8%) | -0.212 (+87.5%) | -37883.3 (+11.0%) | -37334.2 (+25.9%) | -36737.6 (+7.0%) | -37835.3 (+72.5%) |
| Avg. Imp. | +11.8% | +18.3% | +8.5% | +71.3% | +3.6% | +5.1% | +2.0% | +49.4% |
| Power | Congestion | |||||||
| 1st + GAP-LA | 2nd + GAP-LA | 3rd + GAP-LA | [33] + GAP-LA | 1st + GAP-LA | 2nd + GAP-LA | 3rd + GAP-LA | [33] + GAP-LA | |
| case1 | 0.646078 (+0.0%) | 0.646074 (+0.0%) | 0.646078 (-0.0%) | 0.646082 (+0.0%) | 6655094 (-13.8%) | 6615019 (-12.1%) | 6684393 (+16.5%) | 6668336 (-14.7%) |
| case2 | 3.05347 (+0.0%) | 3.05335 (-0.0%) | 3.05335 (+0.0%) | 3.05345 (-0.0%) | 22926581 (-7.5%) | 22565487 (-10.2%) | 22404691 (+11.8%) | 24564053 (-7.9%) |
| case3 | 2.93845 (+0.1%) | 2.93804 (+0.1%) | 2.93750 (+0.0%) | 2.93792 (+1.4%) | 13512436 (-1.5%) | 14294566 (-4.4%) | 13539305 (+13.7%) | 13774919 (-6.2%) |
| case4 | 0.143804 (-0.5%) | 0.143590 (+1.1%) | 0.143594 (+1.1%) | 0.143795 (+1.2%) | 2691325 (+10.9%) | 2667440 (-7.6%) | 2691663 (+22.4%) | 2695234 (-36.7%) |
| case5 | 7.58974 (+1.3%) | 7.57485 (+2.1%) | 7.55822 (-0.1%) | 7.57887 (+3.2%) | 67951941 (-23.5%) | 68592745 (-39.8%) | 68488719 (+29.8%) | 68206894 (-54.6%) |
| case6 | 23.1439 (+1.3%) | 23.1022 (+2.0%) | 23.0783 (-0.1%) | 23.0946 (+2.6%) | 246164776 (-3.9%) | 245829220 (-28.0%) | 239426580 (+17.4%) | 252417493 (-30.7%) |
| case7 | 0.156168 (-0.0%) | 0.156157 (-0.0%) | 0.156148 (-0.0%) | 0.156158 (-0.0%) | 4382560 (-0.8%) | 4410642 (+5.2%) | 4383521 (+0.1%) | 4417159 (-2.5%) |
| case8 | 0.304879 (-0.1%) | 0.304863 (-0.0%) | 0.304885 (-0.0%) | 0.304852 (-0.0%) | 21601536 (+4.3%) | 22622941 (-6.2%) | 21556732 (-4.8%) | 24942228 (+5.5%) |
| case9 | 0.136300 (-0.0%) | 0.136306 (-0.0%) | 0.136301 (-0.0%) | 0.136310 (+0.2%) | 13167451 (+6.4%) | 15576558 (+4.7%) | 13799305 (-1.2%) | 13475394 (-3.1%) |
| case10 | 0.143850 (-0.6%) | 0.143695 (+0.8%) | 0.143565 (+1.1%) | 0.143783 (+1.2%) | 2074929 (+7.0%) | 2077647 (+2.1%) | 2059059 (+5.1%) | 2113216 (-24.6%) |
| case11 | 8.19537 (+0.9%) | 8.14163 (+1.7%) | 8.15759 (+0.1%) | 8.17356 (+5.3%) | 68079290 (-29.0%) | 69214810 (-39.2%) | 68307458 (+48.3%) | 68304067 (-60.2%) |
| case12 | 24.0157 (+1.1%) | 23.9685 (+1.8%) | 23.9422 (-0.1%) | 23.9641 (+2.2%) | 235753772 (+2.9%) | 241013713 (-12.6%) | 236685305 (+29.4%) | 240225253 (-22.3%) |
| Avg. Imp. | +1.1% | +1.8% | -0.1% | +2.6% | -4.6% | -20.9% | +26.1% | -27.7% |
The numbers in the brackets denote the improvement w.r.t the original results of 1st, 2nd, 3rd winners and [33], respectively, by replacing their layer assignment with GAP-LA.
Positive numbers indicate improvement in metrics and negative numbers indicate a degradation.
@@
We implemented our GPU-accelerated layer assignment algorithm using C++ and CUDA on top of HeLEM-GR [33] and HeteroSTA [5, 7, 9, 6], and evaluated the results on benchmarks from the ISPD25 contest [15] released by NVIDIA. The ISPD25 contest released 12 industrial designs synthesized with the modified Nangate 45nm open library technology node and specified an official evaluation flow to assess routing solution performance, as detailed in Section II. The details of 12 designs and the Nangate 45nm library are shown in Table II and I, respectively. Note that the modified Nangate 45nm library holds smaller resistance and capacitance on high layers. All experiments were conducted on a Linux server equipped with a 32-core Intel Xeon Platinum 8358 CPU@2.60GHz and an NVIDIA A800 GPU.
To match the contest evaluation setup, we used the official evaluator for overflow assessment and OpenROAD for evaluating TNS, WNS, and power consumption utilizing 8 CPU threads.
@@ IV-A Main Comparison with Top-3 Winners in the Contest
We obtained the top-3 contest winners’ results and reproduced their experiments on our hardware for comparison. Table III presents a comparison of quality score, runtime and via count against the ISPD 2025 contest’s top-3 winners. Our framework achieved the highest quality score in 9 out of 12 benchmark cases, with comparable performance in the remaining 3 cases. While our static timing analysis engine incurs a 25% runtime overhead due to parsing netlist (.v) and Synopsys design constraints (.sdc) files, the optimization method remains efficient and effective, demonstrating competitive performance relative to the top-3 winners. Our GPU-accelerated layer assignment method shows over 2.18 speedup on average, and over 2.6 speedup on large cases over the multi-threaded version on CPU. As Table III shows that high layers like metal9 and metal10 have significant lower resistance and capacitance, more via usage implies more radical optimization towards timing. The 2nd winner tries to use the fewest resources on high layers and the 3rd winner tries to use the most resources, while our framework follows a more balanced way.
@@
Table IV presents the WNS, TNS, power consumption, and congestion results of four routers across 12 benchmark cases. The 2nd-place winner produces solutions with favorable performance and low congestion, while the 3rd-place winner demonstrates better performance at the expense of higher congestion. The 1st-place winner achieves a more balanced trade-off among metrics [15]. Our method delivers superior performance with acceptable congestion increase, which can be attributed to the critical net analysis-based net ordering and batching strategy. As shown in Table IV, compared to the 1st-place winner, our router achieves 9.9% better WNS, 4.5% better TNS, 1.1% lower power consumption, and 3.8% higher congestion. Versus the 2nd-place winner, we achieve 0.3% better WNS, 5.4% better TNS, 1.5% lower power consumption, and 15.5% higher congestion. In contrast, the 3rd-place winner shows only 0.4% lower power consumption but suffers 4.4% worse WNS, 2.0% worse TNS, and 35.2% higher congestion compared to our approach.
IV-B Comparison with Top-3 Winners and Congestion-driven Router Using Our Method
To better illustrate the effectiveness of our proposed method, we compress the 3D routing solution given by the top-3 winners into 2D net topology and apply our method to them. Table V shows the comparison of overall quality score, WNS, TNS, power and congestion with their original flow. The runtime comparison is omitted here since we only have the binaries and cannot extract their layer assignment runtime. As the contest targets at optimizing the overall quality scores, using GAP-LA on top of the 2D net topology from the top-3 winners can lead to 0.501, 0.933, and 0.817 better quality scores, respectively. In more detail, our method gains 11.8%, 18.3%, 8.5% better WNS, 3.6%, 5.1%, 2.0% better TNS than their original flow, respectively. Although we sacrifice 4.6% and 20.9% congestion compared with the original flow of 1st and 2nd winner, we improve the performance a lot, given the affect that the quality scores from the contest emphasize more on timing metrics. Compared with the 3rd winner, our method can achieve better performance with less congestion. As for the congestion-driven layer assignment method [33], our method achieves 71.3% better WNS, 49.4% better TNS and 2.6% less power with 27.7% more congestion on average, which emphasizes the importance of leveraging performance-aware strategy on our benchmarks.
IV-C Comparison with timing-aware method [23]
There are several works focusing on timing-aware layer assignment [32, 11, 23], among which [23] is the most recent one. To detail our advantages, we compare WNS, TNS, power and congestion with [23] on three representative cases in Table VI. We sincerely appreciate the help from the authors of [23] to run the evaluation and obtain the permission to show these results. However, the comparison is definitely unfair due to design imcompatibility and different optimization objectives. The objective of [23] is to minimize the timing cost under the congestion constraint instead of optimizing the slack with acceptable overflow in our problem setting. Therefore, the comparison is just for reference and does not downplay the effectiveness of [23].
IV-D Ablation Studies
| case1 | case2 | case3 | case4 | case5 | case6 | case7 | case8 | case9 | case10 | case11 | case12 | Avg. | ||
@@ IV-D Ablation Studies
Table VII shows that our framework achieves better WNS results in 11 of the 12 ISPD2025 benchmark cases and better TNS results in all 12 cases. The net ordering and batching strategy enables more precise optimization of critical nets, significantly outperforming the naive strategy that ignores timing analysis information. This advantage is particularly evident in cases sensitive to critical net delay and capacitance, such as case 9. The look-ahead strategy also proves crucial in certain cases, including case 4.
Table VIII presents the number of critical paths and nets for several selected cases. The results show that only a small fraction of nets (less than 1% of all nets) in the netlist play a significant role in timing closure. In case 8, focusing optimization efforts on the 9 critical nets with high timing cost weights is sufficient to achieve zero slack, demonstrating the effectiveness of our critical net selection strategy.
@@
In this paper, we present a GPU-accelerated performance-driven layer assignment framework that generates high-quality routing solutions from 2D routing results. The framework integrates a static timing analysis engine to schedule optimal net ordering and task batching for effective performance optimization. To address the delay cost optimization of large-scale designs with up to 10 million nets, we design GPU kernels for layer assignment using a look-ahead strategy. Compared with the top-3 winners of the contest, our router achieves 0.3%–9.9% better WNS and 2.0%–5.4% better TNS on average with acceptable runtime overhead. We apply our layer assignment framework to the 2D net topology of the top-3 winners and congestion-driven router [33]. Compared with the original flow of top-3 winners, our method gains 0.501-0.933 better quality score, 8.5%-18.3% better WNS, 2.0%-5.1% better TNS on average, respectively.
VI Acknowledgments
Sincere thanks are given to Prof. Genggeng Liu and his student Zepeng Li for their help to run the evaluation of [23] and permission to show these results.